You are going to invest $25,000, part at 14% and part at 16%. What is the most that can be invested at 14% to make at least $3,600 interest per year?
Answer by mananth(16946)

Investment I 14.00% per annum
Investment II 16.00% per annum

x + y 25000 ------------------------1
14.00% x + 16.00% y = $3,600.00
Multiply by 100
14 x + 16 y = $360,000.00 --------2

Multiply (1) by -14
we get

-14 x -14 y = -350000.00

Add this to (2)

0 x 2 y = $10,000.00

divide by 2

y = $5,000.00 investment at 16.00%
Balance $20,000.00 investment at 14.00%
CHECK
$20,000.00 $2,800.00
$5,000.00 $800.00
Total $3,600.00